# Model Informed Flows for Bayesian Inference of Probabilistic Programs

### v1

@@ D.2 Implementation Details
For affine flows, such as those used in the ablation studies in Section 4.3, each conditioning network (for $m_{i},\log s_{i},$ or $t_{i}$) consists of a single linear layer. This layer directly maps the conditioning inputs to the required output parameters.
For flows requiring greater representational capacity, specifically in experiments evaluating network expressiveness (Figure 2) and for the comprehensive benchmarks (Table 3), the conditioning networks are implemented using a two-component MLP-based architecture. Each parameter ($m_{i},\log s_{i},$ or $t_{i}$) is computed by summing the outputs of two parallel components, both of which operate on the same input conditioning variables:
@@
    An MLP component: This is a multi-layer perceptron (MLP) featuring a single hidden layer with ReLU activation functions, followed by a linear output layer. The size of the hidden layer (number of hidden units) is varied in experiments that assess the impact of network capacity (e.g., from $2^{0}$ to $2^{10}$ as illustrated in Figure 2).
@@
For experiments involving baseline models from Blessing et al. [10], we utilized the implementations provided in their publicly available code repository to ensure consistency and fair comparison.

### v2

@@ D.2 Implementation Details
For affine flows, such as those used in the ablation studies in Section˜5.3, each conditioning network (for $m_{i},\log s_{i},$ or $t_{i}$) consists of a single linear layer. This layer directly maps the conditioning inputs to the required output parameters.
For flows requiring greater representational capacity, specifically in experiments evaluating network expressiveness (Figure˜2) and for the comprehensive benchmarks (Table˜3), the conditioning networks are implemented using a two-component MLP-based architecture. Each parameter ($m_{i},\log s_{i},$ or $t_{i}$) is computed by summing the outputs of two parallel components, both of which operate on the same input conditioning variables:
@@
    An MLP component: This is a multi-layer perceptron (MLP) featuring a single hidden layer with ReLU activation functions, followed by a linear output layer. The size of the hidden layer (number of hidden units) is varied in experiments that assess the impact of network capacity (e.g., from $2^{0}$ to $2^{10}$ as illustrated in Figure˜2).
@@
For experiments involving baseline models from Blessing et al. [blessing2024beyond], we utilized the implementations provided in their publicly available code repository to ensure consistency and fair comparison.
D.3 Training Time Comparisons
We report wall-clock training times (in seconds) across six benchmarks and varying hidden sizes. Table˜6 compares our default Model-Informed Flow (MIF) against a variant that conditions on the base noise (“MIF with $\epsilon$”), while Table˜7 ablates the number of conditioning subnetworks: “3 networks” uses $(m,s,t)$ and “2 networks (no tt)” uses only $(m,s)$. Overall, “MIF with $\epsilon$” is consistently faster than the default MIF at comparable capacities, and removing the translation network ($t$) yields additional speedups with similar trends across datasets. We use $D$ to denote latent dimensionality.
Table 6: Training time comparisons (seconds) between the default Model-Informed Flow (MIF) and a variant that conditions on the base noise (“MIF with $\epsilon$”).
Model D Hidden Units MIF (s) MIF with $\epsilon$ (s)
Eight Schools 10 0 $17.06$ $13.88$
$64$ $28.40$ $23.22$
$256$ $28.04$ $22.18$
$512$ $28.21$ $22.03$
$1024$ $30.30$ $22.61$
Funnel 10 0 $18.23$ $14.70$
$64$ $27.38$ $23.16$
$256$ $27.04$ $22.42$
$512$ $28.57$ $22.30$
$1024$ $30.77$ $22.69$
German Credit 125 0 $63.81$ $16.06$
$64$ $148.73$ $27.81$
$256$ $208.44$ $54.92$
$512$ $275.74$ $93.26$
$1024$ $417.05$ $173.51$
Radon 174 0 $84.25$ $15.78$
$64$ $238.04$ $37.32$
$256$ $331.24$ $94.82$
$512$ $453.70$ $168.28$
$1024$ $699.73$ $338.63$
IRT 143 0 $72.26$ $17.59$
$64$ $196.64$ $33.15$
$256$ $253.17$ $67.95$
$512$ $457.56$ $119.95$
$1024$ $516.98$ $222.65$
MovieLens 882 0 $458.28$ $28.97$
$64$ $1981.14$ $384.45$
$256$ $4847.23$ $1387.64$
$512$ $8634.91$ $3156.82$
$1024$ 15 873.4515\,873.45 $6749.33$
Table 7: Ablation on the number of conditioning subnetworks (seconds). “3 networks” uses shift, scale, and translation $(m,s,t)$. “2 networks (no tt)” removes the translation network and uses only $(m,s)$ under the same training setup.
Model D Hidden Units 3 Networks (s) 2 Networks (no $t$) (s)
Eight Schools 10 0 $17.06$ $16.98$
$64$ $28.40$ $24.13$
$256$ $28.04$ $24.81$
$512$ $28.21$ $25.80$
$1024$ $30.30$ $25.76$
Funnel 10 0 $18.23$ $18.63$
$64$ $27.38$ $25.89$
$256$ $27.04$ $25.85$
$512$ $28.57$ $26.27$
$1024$ $30.77$ $26.01$
German Credit 125 0 $63.81$ $52.68$
$64$ $148.73$ $115.39$
$256$ $208.44$ $139.90$
$512$ $275.74$ $175.94$
$1024$ $417.05$ $258.98$
Radon 174 0 $84.25$ $77.52$
$64$ $238.04$ $220.17$
$256$ $331.24$ $306.44$
$512$ $453.70$ $419.42$
$1024$ $699.73$ $647.75$
IRT 143 0 $72.26$ $66.73$
$64$ $196.64$ $181.92$
$256$ $253.17$ $234.43$
$512$ $457.56$ $423.00$
$1024$ $516.98$ $478.45$
MovieLens 882 0 $458.28$ $423.65$
$64$ $1981.14$ $1832.85$
$256$ $4847.23$ $4486.69$
$512$ $8634.91$ $7992.34$
$1024$ 15 873.4515\,873.45 14 687.9914\,687.99